### v1

# Generation and Preservation of Large Entangled States on Physical Quantum Devices

Abstract
As quantum technology advances and the size of quantum computers grow, it becomes increasingly important to understand the extent of quality in the devices. As large-scale entanglement is a quantum resource crucial for achieving quantum advantage, the challenge in its generation makes it a valuable benchmark for measuring the performance of universal quantum devices. In this work, we study entanglement in Greenberger-Horne-Zeilinger (GHZ) and graph states prepared on the range of IBM Quantum devices. We generate GHZ states and investigate their coherence times with respect to state size and dynamical decoupling techniques. A GHZ fidelity of is measured on a 32-qubit GHZ state, certifying its genuine multipartite entanglement (GME). We show a substantial improvement in GHZ decoherence rates for a 7-qubit GHZ state after implementing dynamical decoupling, and observe a linear trend in the decoherence rate of µs for up to qubits, confirming the absence of superdecoherence. Additionally, we prepare and characterise fully bipartite entangled native graph states on 22 IBM Quantum devices with qubit counts as high as 414 qubits, all active qubits of the 433-qubit Osprey device. Analysis of the decay of 2-qubit entanglement within the prepared states shows suppression of coherent noise signals with the implementation of dynamical decoupling techniques. Additionally, we observe that the entanglement in some qubit pairs oscillates over time, which is likely caused by residual ZZ-interactions. Characterising entanglement in native graph states, along with detecting entanglement oscillations, can be an effective approach to low-level device benchmarking that encapsulates 2-qubit error rates along with additional sources of noise, with possible applications to quantum circuit compilation. We develop several tools to automate the preparation and entanglement characterisation of GHZ and graph states. In particular, a method to characterise graph state bipartite entanglement using just 36 circuits, constant with respect to the number of qubits.
I Introduction
Producing large-scale entangled states on many qubits represents an important frontier in the race to scale up physical quantum computers. In quantum computers, entanglement manifests as non-classical correlations between qubits, such that qubits involved in an entangled system cannot be described independently from each other [1, 2]. Entanglement has been found to play a significant role in achieving quantum advantage [3, 4, 5, 6], and many quantum information processing protocols explicitly rely on entanglement as a resource [7, 8, 9, 10]. Furthermore, entanglement over large arrays of qubits is essential in many fault-tolerant computation schemes [11, 12, 13]. Multipartite entanglement is known for its complex structure [1], and modern quantum devices have passed the number of qubits a classical computer can store an arbitrary quantum state on [14, 15]. Consequently, as noisy intermediate-scale quantum (NISQ) devices [16] continue to improve in both size and error rates, it becomes paramount to characterise the ability of a quantum computer to generate and maintain large entangled states in a scalable manner.
Verifying multipartite entanglement on a quantum device requires measuring the fidelity of a highly entangled multi-qubit state. Greenberger-Horne-Zeilinger (GHZ) states [17] are a popular choice, as measuring a GHZ state fidelity of greater than 0.5 is sufficient for verifying genuine multipartite entanglement (GME). For quantum devices with full qubit control, GHZ states of sizes 27 and 29 qubits have been observed on superconducting systems [18, 19], and GHZ states of size 32 qubits have been observed on ion-trap systems [20]. Graph states, also known as cluster states, are another widely studied class of multipartite entangled states. Graph states are useful for showing mixed state bipartite entanglement, and full bipartite entanglement has been observed on up to 65-qubit graph states in superconducting systems [21, 22, 23], and 20-qubit graph states in ion-trap systems [24]. There has also been recent work showing violation of robust Bell inequalities for 57-qubit path graph states [25], and genuine entanglement for 51-qubit path graph states and 30-qubit 2D graph states [26] on superconducting devices.
Increasing the size of multipartite entanglement is profitless if the entangled state degrades too rapidly to be able to perform meaningful operations on. A source of concern when engineering physical quantum devices is superdecoherence; a phenomenon where qubit decoherence rates scale linearly with the number of qubits due to the coupling of multiple qubits with a single reservoir [27]. Studying the decay of entangled states can reveal information about the noise underlying a quantum system. For example, the decoherence of GHZ states has been used to show superdecoherence (or the lack thereof) in superconducting and ion-trap systems [28, 29].
Techniques such as dynamical decoupling have been developed to suppress the effect of environmental noise on quantum states. Dynamical decoupling, a quantum control technique that employs sequences of control pulses to suppress the effect of environmental noise on quantum states, has been shown to be remarkably effective at protecting four-qubit GHZ states on ion-trap devices [30].
In this work, we investigate both GME in the form of GHZ state fidelities and mixed bipartite whole-device entanglement using native-graph states. We develop several tools to automate the preparation and verification of these states over the range of IBM Quantum devices. In particular, we employ an automated GHZ state embedding scheme that embeds tree-type GHZ preparation circuits with minimum-depth on heavy-hexagonal qubit architectures to generate up to 32-qubit GHZ states on the 127-qubit ibm_washington device. Using multiple quantum coherences (MQC) [31], we record a fidelity of for the 32-qubit state after mitigating for measurement errors.
[…]
We finally investigate dynamical decoupling for preserving native-graph state entanglement. Notably, we observe revivals in entanglement signals for several qubit pairs. The observed resurgent signals in negativity are consistent with signals produced by residual ZZ interactions, which are known to affect superconducting transmon qubits and generate local entanglement [32, 33, 34, 35, 36]. We observe improvement in mean device entanglement lifetimes after implementing PDD.
II Generation and Decay of Greenberger-Horne-Zeilinger States
GHZ states and verifying genuine multipartite entanglement
GHZ states [17] are highly entangled multipartite states that are uniquely fragile to noise, where a single-qubit phase error can destroy the whole-state entanglement. The ability of a quantum device to generate large GHZ states with high fidelity depends on a holistic combination of factors, including qubit count, gate error rates, coherence times, and cross-talk. Generating and verifying such states has therefore become a valuable benchmark for NISQ devices. GHZ states are typically prepared by initializing a source qubit in the state, and then iteratively applying CNOT gates from the source qubit (or any other qubit that has already had a CNOT applied in this manner) to all other qubits prepared in the state (Fig. 1). The resulting state is an equal superposition of all subsystems in the state and all subsystems in the state. Formally, an -qubit GHZ state is expressed as:
[…]
To certify -qubit GME, it is sufficient to show a GHZ fidelity of greater than 0.5 [37] (note that a GHZ state may be GME with a fidelity under 0.5). The fidelity between a pure target state and the actual (noisy) state is calculated as
The resource requirement for full quantum state tomography scales exponentially with the number of qubits, making it intractable to obtain using this approach even for medium-sized systems. Fortunately, GHZ states of all sizes have the convenient property that their density matrices (ideally) consist of only four non-zero corner elements. As a result, for GHZ states, Eq. 2 can be re-expressed as
where are the directly measured populations, and are the coherences which can be measured using either parity oscillations [37, 28] or multiple quantum coherences (MQC) [31].
MQC is advantageous due to its natural integration with dynamical decoupling-based techniques, such as the Hahn echo, which refocuses noise and mitigates dephasing, as well as readout error mitigation, since readout errors are typically the dominant noise factor for low-depth circuits. MQC has been used to verify GHZ states of up to 29 qubits on superconducting quantum devices [31, 18, 19]. The methodology for computing GHZ coherences via MQC can be summarized as follows:
1.
Prepare the -qubit GHZ state in the form as exemplified in Fig. 1b.
2.
(Optionally) Apply a refocusing -pulse (i.e. an -gate) on each qubit, producing the state .
3.
Apply a phase rotation of on each qubit in the GHZ state, adding an accumulative phase of to the overall state: .
4.
Disentangle the state by applying the inverse of the GHZ state preparation circuit from step 1. The accumulated phase is mapped onto qubit 0: .
5.
Obtain the overlap signal as the probabilities of measuring the state for different values of .
The GHZ coherence can then be calculated as where are the overlap signal amplitudes which can be obtained by applying a Fourier transform to :
where is the number of angles in the summation. To detect up to frequency , we measure for angles , . Further details including deriving the fidelity from the overlap signal concretely is presented in [31, 18].
GHZ state embedding on physical topologies
When preparing GHZ states on physical devices, accounting for hardware topology and gate error rates is crucial for maximizing the final state fidelity. In previous experiments [18], circuit embedding was performed manually. While suitable for smaller devices, the introduction of 433-qubit processors, and most recently, the announcement of a 1121 qubit processor [38] necessitate methods for automated embedding of device entangling circuits.
In this work, we develop a topology-agnostic GHZ state preparation scheme that constructs tree-type GHZ preparation circuits on heavy-hexagonal layouts with minimal depth. The method additionally involves selecting qubit subsets with optimized parameters such as low two-qubit gate error rates. The exact parameter or combination of parameters to optimise is specified by assigning corresponding weights to the edges of the graph. We initially considered only the CNOT error rate. Our algorithm is divided into two components. The first component embeds a GHZ circuit as a directed tree branching out from a single source qubit. We apply CNOT gates in parallel prioritizing least depth first and lowest two-qubit error rate second. The second component runs the first component algorithm multiple times, trialling each physical qubit or a subset of physical qubits as the source qubit. We then select the circuit with the least depth as the primary criterion and the lowest total cost parameter as the secondary criterion. This way, we can conveniently embed least-depth GHZ states on larger and larger quantum devices.
IBM employs a heavy-hexagonal lattice as its principal architecture for all their devices. The choice in topology is motivated by a reduction in qubit frequency collisions [39] and spectator errors [40], as well as surface code versatility [41]. It has been shown for star graph states (which are equivalent to GHZ states under LOCC) that the CNOT depth to construct states of size scales as approximately on heavy-hex architectures [25]. Precisely, a GHZ circuit embedded on an infinite heavy-hexagonal lattice with depth can prepare states of up to size
where is the number of qubits.
For the heavy-hex topology, our algorithm embeds GHZ states with this optimal depth scaling up to a limit imposed by the boundaries of the physical device. We illustrate this for a 127-qubit Eagle processor in Fig. 2, where it can be seen that a tree-type GHZ state with a centrally located source qubit closely follows Eq. 5, whereas selecting source qubits closer to the boundaries results in worse scaling. We showcase an algorithmically embedded example 60-qubit GHZ state on the same processor in Fig. 3. We note that the algorithm is designed to be compatible with any finite-degree device graph, although its performance on other topologies has not yet been investigated.
Scalable quantum readout-error mitigation (M3)
Readout or measurement errors represent the largest source of noise for low-depth circuits executed on NISQ devices. Readout error rates of even a few percent can be debilitating to the output fidelity of otherwise well-performing systems. Nevertheless, if the readout error takes on a predominantly classical form, which has been shown to be largely true for IBM Quantum transmon devices [18], its effects on measured probability distributions can be mitigated via post-processing. In its simplest form, quantum readout error mitigation (QREM) solves the linear equation
where is the noisy probability vector returned by the system, is the ideal probability vector in the absence of readout errors (but may include other errors), and is the calibration matrix, where corresponds to the probability of bit string being measured as bit string .
There are limitations to QREM, with the most significant one being the exponential scaling of classical resources required to solve for with respect to . There are a variety of approaches to overcome this challenge, which often involve approximating the calibration matrix by reducing it to a tensor product of single-qubit components [18, 42]. More recently, a measurement error mitigation method called M3 (Matrix-Free Measurement Error Mitigation) has demonstrated order-of-magnitude improvement in mitigation time over traditional methods [42]. M3 relies on two main optimizations: subspace reduction of the full calibration matrix based on noisy input bit strings, and implementation of a matrix-free, iterative method for solving the system of linear equations. The method involves directly computing matrix elements as
This error mitigation technique has some caveats. Firstly, M3 works natively with quasi-probability distributions which can contain negative elements. These non-physical probabilities arise from the finite sampling and, while still adding up to one, are incompatible with methods such as MQC. Therefore, a classical algorithm is used to efficiently convert the quasi-probabilities into the closest physical probability distribution under the -norm, and it runs in time [43]. Furthermore, the speedup of the mitigation process depends on the sparsity of the measured probability distributions. A sparser corresponds to a greater subspace reduction, which makes M3 optimal for GHZ states since they have only two measurement outcomes. However, scalable mitigation of readout errors comes at the expense of increased uncertainty for measurement outcomes. The mitigation overhead is given by:
where is the trace norm of . Quantity gives an upper bound to the standard deviation of an observable
where is the number of samples. Thus, results mitigated using M3 will require more samples to achieve similar uncertainty with results mitigated using traditional QREM.
Nevertheless, the benefits of employing M3 heavily outweigh the limitations in this use case. For our GHZ experiments, we employ M3 using the publicly available mthree Python package. We apply the correction algorithm with all the default settings, which include correcting bit strings up to a Hamming distance equal to GHZ size , a convergence tolerance of the iterative method of , and a maximum number of iterations of 25.
Verifying 32-qubit GHZ states on a 127-qubit Eagle Processor
We prepare and measure the fidelities of tree-type GHZ states of sizes , on the 127-qubit ibm_washington device. GHZ state preparation circuits are constructed using the embedding algorithm. The algorithm selects qubit 73 as the source qubit, although this choice varied across calibration cycles due to physical error drift. Notably, all GHZ sizes are embedded with an optimal CNOT depth of 8, which is not possible on any of IBM Quantum’s smaller heavy-hex devices due to boundary effects. We conduct five sets of experiments to obtain five independent measurements of the GHZ fidelity for each . Each experiment requires circuits since the MQC method measures the overlap signal for values of (plus one circuit to measure the population). Prior to any GHZ experiments, we perform M3 readout error calibration. All circuits are executed with 4196 shots each.
Figure 4 shows population, coherence, and fidelity plots for GHZ states of size . Plotted data points display the mean value across five experiments and error bars represent the standard error. We plot results with and without readout error mitigation using M3. After applying readout correction, we measure the fidelity lower bounds for all states to be above the 0.5 threshold required to demonstrate GME. In particular, we measure a mitigated fidelity of for the 32-qubit GHZ state. To the best of our knowledge, this is the largest GHZ state observed to have a fidelity of over 0.5. Furthermore, Fig. 5 shows the corresponding MQC overlap signals (with QREM applied). The signal amplitudes are filtered and calculated using a fast fourier transform algorithm. The observed phase shift for certain is likely caused by free rotations in idle qubits.
The results are unusual in some aspects. Most obvious are the anomalously high fidelities for the 30-qubit GHZ state, where we measure a mitigated fidelity of . Additionally, the measured populations seem to increase for GHZ states of size , especially after M3 is applied. These peculiarities are likely due to several factors. Firstly, the experiments were executed in reverse order where the 32-qubit GHZ experiments were performed first. Due to the nature of the IBM Quantum job queuing system, not all experiments could be executed consecutively. In fact, experiments for GHZ size were performed in a different calibration cycle than experiments for GHZ size . This is important because over the course of our research, we observed considerable performance drift in the ibm_washington device. In some attempts, the measured GHZ fidelities were well below the 0.5 threshold with no changes to the experimental procedure. We suspect that the discontinuity in the results reflects a decline in device performance during the latter half of experiments.
We draw comparisons to previous GHZ experiments on IBM Quantum devices [31, 18, 19]. We observe a relatively small decrease in measured fidelities with increasing size. This may be explained by all GHZ sizes being prepared by the same CNOT depth of 8, in addition to the previously mentioned performance drift. We remark that the measured fidelities for similar size GHZ states on ibm_washington are not substantially higher than the GHZ fidelities on smaller devices from previous experiments [18, 19] (in some cases being lower). This is not unexpected, since the average device error rates are often lower on the largest devices. We postulate that the larger verifiable GHZ state sizes on the 127-qubit device are partially enabled by its sheer scale—which allows larger GHZ states to be prepared with lower circuit depth. This highlights the importance of the scale of a quantum device in addition to the quality of its qubits. Next, we study the decay of GHZ states over time.
III Preserving GHZ States via Dynamical Decoupling
@C=0.8em @R=0.7em
[…]
@C=0.8em @R=0.5em
[…]
@C=0.8em @R=0.85em
[…]
In this section, we explore the potential of dynamical decoupling techniques in prolonging the lifetimes of GHZ states on IBM Quantum devices. Dynamical decoupling is an open-loop control technique that mitigates decoherence in quantum computers by implementing sequences of control pulses [44, 45]. In theory, these control sequences effectively average out undesirable couplings between qubits and their environment. Dynamical decoupling can be seen as a generalization of the Hahn spin echo [46], which is the special case for a single or pair of control pulses.
There are variations between dynamical decoupling schemes. The most basic scheme, known as periodic dynamical decoupling (PDD), applies equally spaced control pulses in quick succession. More advanced schemes include bounded-strength continuous sequences [47], concatenated dynamical decoupling (CDD) [48], and Uhrig dynamical decoupling (UDD) [49]. Different dynamical decoupling schemes may optimize for different scenarios and noise environments. There is ongoing research investigating the best way to integrate dynamical decoupling protocols with quantum computing algorithms.
A typical dynamical decoupling technique on IBM Quantum devices is to implement control sequences during qubit idle periods. Most relevantly, dynamical decoupling has been shown to preserve GHZ coherence times by orders of magnitude on ion-trap qubits for up to four-qubit GHZ states [30]. We study the efficacy of integrating dynamical decoupling with MQC in order to preserve GHZ fidelities. The experimental procedure follows from previous sections with the exception of adding a variable delay period between state preparation and fidelity measurements. We compare the decay of GHZ populations, coherences, and fidelities for free decays (idle qubits), and decays preserved with Hahn echo and PDD. The relevant circuit diagrams are shown in Fig. 6.
Hahn Echo and Periodic Dynamical Decoupling for 7-Qubit GHZ States
Decay experiments are performed for 7-qubit GHZ states prepared on the ibmq_mumbai device consisting of 27 qubits and a quantum volume of 128. For free decays, we increment the circuit delay by 1 µs up to a maximum delay of 15 µs. For decays preserved with Hahn echo or PDD, we increment by 2 µs up to a maximum of 30 µs. For PDD, we implement control-pulses in 0.5 µs regular intervals. We obtain both unmitigated results and results with QREM-applied via the M3 protocol. The GHZ state populations, coherences and fidelities are measured across five independent experiments, where circuits are executed with 4196 shots each.
Figure 7 displays the experimental results. As shown in Fig. 7a, neither Hahn echo nor PDD led to a marked improvement of GHZ population times. In fact, PDD appears to accelerate the decay of GHZ populations. The results are related to how the ground and excited state populations evolve with respect to relaxation errors on superconducting quantum devices. In detail, qubits in the excited state will eventually spontaneously decay into the ground state at a rate described by the relaxation time. After a sufficient amount of time, a quantum computer will reset to the all ground state (with some fluctuations). The decay in GHZ populations is primarily caused by bit flips in the state due to thermal relaxation, although environmental noise can also cause random bitflips. The application of control pulses in Hahn echo or PDD, which flip the ground and excited state probabilities, will do little to prevent relaxation errors. In fact, as shown in the PDD curve, repeated application of -gates only introduces additional noise from single-qubit gate errors.
Figure 7b shows substantial improvement in GHZ coherence times in experiments with Hahn echo and PDD. The GHZ coherences quantify the non-classical correlations between the and states. For free decays, this correlation drops to effectively zero by around µs. In contrast, the PDD curve maintains a measured GHZ coherence of after µs. Hahn echo also appreciably prolongs GHZ coherence times, resulting in at µs, albeit to a lesser extent. These results highlight the efficacy of dynamical decoupling-based techniques in protecting GHZ states against dephasing errors (related to dephasing times) on superconducting quantum computers.
The decay of GHZ state populations and coherences are combined in Fig. 7c, which plots the GHZ state fidelities as a function of circuit delay . Both Hahn echo and PDD are shown to be effective techniques for preserving GHZ fidelities. It is interesting to observe that although PDD is superior to Hahn echo for preserving GHZ coherences, it is worse at preserving GHZ populations. This is likely due to additional noise introduced from the PDD gate sequences, which has the effect of obfuscating population measurements in exchange for better protection against decoherence. As a result, Hahn echo exhibits roughly similar performance to PDD in preserving overall fidelities. In future experiments, it may be worth testing more advanced dynamical decoupling protocols such as UDD [49], which are shown to be more typically more effective than PDD [50]. We comment that the small differences in the initial values of , and are likely attributed to device drift. We also remark that applying QREM mainly increases the initial , and values with little influence to the decay rates. Next, we evaluate the scaling of GHZ decoherence rates as a function of state size.
Scaling of Decoherence Rates and GHZ Size
Studying the decoherence rates of multipartite states as a function of their size may provide vital insight into the noise underlying a quantum system. The strength and nature of this noise can determine the feasibility of scaling up a quantum device. In particular, it may reveal whether a system exhibits superdecoherence. Superdecoherence describes the coupling of qubits to a single reservoir, which cause qubit decoherence rates to scale with the size of the system [27]. Such an effect is detrimental to the realisation of large-scale, fault-tolerant quantum computers. GHZ states are particularly convenient for detecting superdecoherence due to their high sensitivity to noise. In detail, GHZ states accumulate decoherence between qubits, so if the dominant noise model is uncorrelated across qubits (i.e. the decoherence rate per qubit is constant), we expect GHZ decoherence rates to scale linearly with the number of qubits. In contrast, if the dominant noise model is correlated across qubits (i.e. the system experiences superdecoherence), we expect GHZ decoherence rates to scale polynomially.
[…]
GHZ decoherence rates as a function of state size have been studied on ion-trap and superconducting quantum devices for up to state size 6 and 8, respectively [28, 29]. Prominently, the ion-trap device was shown to exhibit quadratic scaling of GHZ decoherence rates, indicating superdecoherence (note that this does not imply the same for all ion-trap systems). On the other hand, the IBM Quantum superconducting device displayed linear GHZ scaling. In this section, we extend the study on GHZ decoherence scaling on IBM’s more recent superconducting devices, for GHZ states of up to 15 qubits in size. Furthermore, for the first time, we incorporate readout error mitigation and measure the coherences via MQC, incorporating Hahn echo.
To provide easy comparison, we conduct our experiments in a manner similar to Ozaeta & McMahon’s previous study involving IBM Quantum devices [29]. The study, undertaken in 2018, measured GHZ decay rates on the now retired 16-qubit ibmq_melbourne device, which employed a square lattice qubit topology. Today, we implement our study on the 27-qubit ibm_hanoi device, which employs a heavy-hex topology. In contrast to Ozaeta & McMahon’s study which measures GHZ coherences using parity oscillations, we employ MQC, incorporating dynamical decoupling-based techniques in the form a single -pulse.
The circuit for measuring GHZ decoherence rates is shown in Fig. 8. To prolong coherence times, we implement a single -pulse in the middle of the delay period. For an -qubit GHZ state, we model the coherence as a function of delay as the exponential decay
where is the initial coherence, and is the decoherence rate, where its reciprocal is the GHZ coherence time.
GHZ Decoherence Scaling Results
Figure 9a plots the -qubit GHZ state coherences (normalized) as a function of circuit delay for on ibm_hanoi, which consists of 27 qubits and a quantum volume of 64. We increase total delay in increments of 2.5 µs. The maximum delay for each experiment ranges from µs for to µs for . For reference, the average CNOT gate time on ibm_hanoi is 385 ns. Data points represent the average measured coherence between five experiments, and error bars represent the standard deviation. We fit the plotted data with the exponential decay curve Eq. 10. We execute circuits with 4196 shots each. We mitigate readout errors using M3, although as shown in the previous section, this has little effect on the decay rates.
Most of the decay curves are modeled well by the exponential decay function. A notable exception, however, is the coherences, which appear to plateau slightly before exhibiting exponential decay. As a result, the decay fit incorrectly projects the initial coherence to be greater than 1. We observe a clear pattern of increasing decay rates with increasing GHZ size. In order to quantify this trend, we plot the decoherence rate and the GHZ size , shown in Fig. 9b. We take the error from the standard deviation, which we derive from the covariance matrix produced by the fitting algorithm. In agreement with Ozaeta & McMahon, we are able to well fit the data with the linear trendline µs with —now extending up to . (With an anomoalous data point at ). Our results support the notion that recent IBM Quantum transmon devices are naturally robust against superdecoherence. We remark that although we only measure coherence times for GHZ states of up to 15 qubits, MQC and dynamical decoupling-based techniques improve the initial values enough that one can feasibly extend the study to larger GHZ states, especially as error rates and coherence times improve.
[] (MQC) | [µs] (MQC) | [µs] [29] | |
1 | - | - | |
2 | - | - | |
3 | |||
4 | - | - | |
5 | |||
6 | - | - | |
7 | |||
8 | - | - | |
9 | - | ||
11 | - | ||
13 | - | ||
15 | - |
We summarise the GHZ coherence times on ibm_hanoi in Table 1. For easy comparison, we include the measured GHZ coherence times from Ozaeta & McMahon’s experiments. For , we report an average increase of coherence times of 133% over previous results. Furthermore, we report longer coherence times on 15-qubit GHZ states than on previous 8-qubit GHZ states of μs and μs, respectively. The improvement in GHZ coherence times may be attributed to hard improvements in combination with superior coherence detection methods, incorporating dynamical-decoupling based techniques.
IV Efficient Bipartite Entanglement Characterisation in Whole-Device Graph States
Graph States
Graph states are a class of entangled multi-qubit states that are defined with respect to a connected graph. They are a generalisation of cluster states [51], and hence form a universal basis for measurement-based computation [52]. Graph states are additionally useful for quantum error correcting codes [53], quantum secure communication [54], quantum metrology [55], and probing Bell inequalities over multi-partite systems [56]. In quantum circuit notation, a graph state may be expressed as
where , is the set of edges connecting graph containing vertices (qubits), and represents a controlled-phase gate between adjacent qubits and . We can equivalently define the stabilising operator for each qubit in :
where are Pauli operators acting on qubit and is the set of qubits adjacent to . Thus is the simultaneous eigenstate of operators following .
Graph states are a convenient choice for studying large-scale entanglement as they are simple to prepare and comparitively noise robust [51]. Controlled-phase operations that do not overlap vertices can be applied in parallel, allowing any graph state to be prepared by a linear-size constant-depth circuit [57]. Concretely, a bounded degree graph can be prepared with a two-qubit gate depth equal to the maximum degree between its vertices [58]. An example least-depth graph state preparation circuit on a seven-qubit layout is shown in Fig. 10. The qubit layout is represented in Fig. 10a where nodes represent qubits and edges display possible CNOT operations. Since the graph has a maximum node degree of three, its graph state preparation circuit has a two-qubit gate depth of three.
[…]
To characterise bipartite entanglement across an entire device, we prepare a native-graph state containing every edge. We then perform full quantum state tomography on every locally entangled bipartition corresponding to each edge on the device. In detail, graph states have the property that projecting all but two qubits in an entangled cluster leaves the pair in a maximally entangled Bell state [59]. The extent of two-qubit entanglement can then be quantified by measuring the negativity [60]. For a quantum state represented by the density matrix , the negativity between subsystems and is calculated as
where are the eigenvalues of , where is the partial transpose of with respect to subsystem . A maximally entangled Bell state has a negativity of , whereas a fully separable state has a negativity of . Although there are many other entanglement measures, such as entropy of entanglement and concurrence, negativity is chosen because it is easy to compute and is related to the teleportation distance [60]. Specifically, the minimum teleportation distance achievable with state , acting on , is bound by
A device is said to be whole-device entangled if every qubit is connected to the main graph where edges correspond to qubit pairs with a measured negativity of . This is distinct from saying the qubits are genuinely multipartite entangled (which follows a more strict criteria), but rather, there exist no bipartition of qubits on the device that results in separable states.
Bipartite Entanglement Characterisation Protocol
We develop a protocol to efficiently characterise bipartite entanglement on quantum computers, inspired by experiments from Mooney et al. [22, 23]. Development of the scheme is driven by three main design principles: to devise an entanglement characterisation protocol that is highly automated, scalable, and architecture-independent. We implement the protocol in Python, utilising the Qiskit API to interface with IBM superconducting quantum devices. However, the techniques and procedures in the program are generally applicable. The protocol can be divided into five components:
1.
Native-graph state preparation. Automatically construct a whole-device graph state preparation circuit which entangles every qubit on the device.
2.
Parallel Quantum State Tomography. Execute quantum state tomography circuits on Bell states prepared on the graph state in parallel. For heavy-hex qubit architectures, this step can be performed in four batches of nine circuits each.
3.
Quantum readout error mitigation. Mitigate readout errors using classical post-processing.
4.
Density matrix reconstruction. Reconstruct the Bell state density matrices using readout error mitigated or unmitigated measurement results.
5.
Negativity calculation and entanglement mapping. Calculate the bipartite negativities corresponding to each qubit pair on the device and construct the entanglement graph.
The first and second component, which contain techniques unique to this work [22, 23], are elaborated below.
IV.0.1 Native-Graph State Preparation
Similar to the GHZ case, the objective is to embed depth-optimal circuits using an automated routine. (In previous works [21, 22, 23], this is done manually) Since heavy-hex lattices have a maximum node degree of three, it is possible to embed a native-graph state circuit with a minimum two-qubit gate depth of three. Topology-specific methods for constructing optimal graph state circuits exist, such as stitching together smaller circuits embedded on unit cells [23]. It is sufficient, however, to implement a greedy algorithm that applies as many CNOT gates in parallel in each step. The algorithm also has the advantage of working with any qubit layout, although it is unknown if it is universally optimal. A whole-device graph state embedding prepared by the algorithm on the 127 qubit Eagle processor is shown in Fig. 11. Additional graph state embeddings on other IBM physical layouts are shown in the results section below.
IV.0.2 Parallel Quantum State Tomography
In previous graph state experiments [22, 23], full quantum state tomography (QST) is performed on one qubit pair (and its neighbours) at a time. Using such method, the number of circuits required to fully characterise a device increases linearly with the number of edges. Our procedure improves upon this by executing QST in parallel. In detail, we perform simultaneous basis measurements on non-overlapping sets of qubits, where each set defines a target qubit pair and its neighbours. By grouping sets into batches and performing parallel QST for batches at a time, we can fully characterise bipartite entanglement on any size device (provided invariable qubit topology) with a constant number of circuits.
Using a specialised scheduling scheme for the heavy hex hardware layout, the number of required batches can be reduced to only four when allowing Bell pairs to share neighbours. This can be achieved by using an alternate tiling of two unit cells (able to be rotated 180°), where each unit cell is composed of six edges and each edge is assigned to one of four batches. However, similar to the case of graph state embedding, we instead opt for a topology-agnostic algorithm that performs parallel QST for as many non-overlapping Bell pairs as it can fit into a single batch. In our implementation, we prevent target pairs from sharing neighbours for practical convenience. Table 2 lists the number of batches required to characterise IBM Quantum devices up to 433 qubits in size.
Device(s) | Qubits | Batches | Circuits |
manila, belem, lima, quito, santiago, bogota | 5 | 4 | 36 |
perth, jakarta, lagos | 7 | 6 | 54 |
guadalupe | 16 | 6 | 54 |
kolkata, mumbai, toronto, montreal, hanoi, cairo | 27 | 6 | 54 |
brooklyn, ithaca | 65 | 6 | 54 |
washington, sherbrooke, brisbane | 127 | 8 | 72 |
seattle | 433 | 8 | 72 |
The table shows that the batching algorithm collates tomography circuits into 6-8 batches for heavy-hex devices. Performing full QST on qubits requires circuits corresponding to each combination of Pauli bases. Therefore each batch, which performs two-qubit QST in parallel, contains nine circuits. The variation in the number of batches is likely attributed to the greedy nature of the algorithm, which may group sets of qubits into batches in non-optimal order. Nevertheless, the procedure’s main utility lies in reducing the number of tomography circuits to a roughly constant number in addition to being compatible with various qubit topologies.
Bipartite Entanglement on IBM Quantum Devices
We characterise bipartite entanglement on all IBM Quantum devices accessible by the University of Melbourne IBM Quantum Hub. At the time of experiment, these include four 5-qubit systems, five 7-qubit systems, one 16-qubit system, eight 27-qubit systems, three 127-qubit systems, and one 433-qubit system, totaling to 22 systems. Both unmitigated and mitigated results using QREM are shown. It is important to present unmitigated results because not all protocols involving graph states can incorporate readout error mitigation. Notably, quantum teleportation schemes which use mid-circuit measurements are incompatible with QREM.
We perform eight sets of graph state experiments per device, sampling all circuits with 8192 shots. Readout error calibration circuits are sampled with the same number of shots. All native-graph states are prepared with the optimal two-qubit gate circuit depth of three. The number of circuits per experiment for various size devices is shown in Table 2. Besides practicality, reducing the number of circuits per experiment is beneficial because it also reduces the variability in results due to device drift. To assign a single negativity for each device edge, we calculate the mean maximum negativity between experiments, where maximum refers to the largest negativity between possible Bell state projections. We take the error to be standard error.
Figure 12 show sample negativity plots for devices up to 27 qubits in size and their respective graph state embeddings. Edges are sorted in order of ascending lower bounds of mitigated negativities. Among these systems, ibm_oslo purports both the highest mitigated and unmitigated mean device negativities of 0.488 and 0.403, respectively. Nevertheless, all negativity plots indicate that each of these systems exhibit whole-device entanglement. The improvement in negativities due to QREM is consistently significant. The average percentage improvement in mean negativity across the four devices is 26.3%
Figure 13 shows the negativity plot for 127-qubit device ibm_washington, with the respective graph state embedding is shown in Fig. 11. The unmitigated negativities have a mean of 0.290 and a standard deviation of 0.117. After implementing QREM, the resulting negativities have a mean of 0.408 and a standard deviation of 0.102—a 40.7% percent improvement in mean device negativity. We comment on several anomalous edges with large gaps between the mitigated and unmitigated negativity, such as edges 4–5, 4–15. These can be attributed to significantly higher than average readout error rates for certain qubits. In particular, qubits 4 and 12 display abnormally high readout error rates of 0.338 and 0.390 respectively, corresponding to the large negativity gaps at edges 4–5, 4–15, 3–4, 12–17, 12–13, and 11-12.
As IBM Quantum’s earliest Eagle processor, ibm_washington purports a lower mean device negativity than most 27-qubit Falcon processors and the newer Eagle r3 devices (see Table 3). Although, after applying QREM, we observe whole-device entanglement across all 127 qubits. To illustrate, Fig. 14 draws graphic representations of entanglement within ibm_washington. Negativity values are mapped on device edges where thin red edges represent low negativity and thick blue edges represent high negativity. Edges with lower-bound negativities of zero are greyed out. We remark that edges with low negativity tend to coalesce in regions. These areas of low entanglement, such as in the lower left corner of Fig. 14, may arise due to physical factors such as non-uniform heat distribution in the device.
To further our investigation of entanglement within the 127-qubit device, we plot the Bell state negativities against CNOT error rates in Fig. 16. Precisely, we take the CNOT error to be the average CNOT error between edges in the tomography set, which in addition to the Bell state pair, includes its adjacent neighbours. Furthermore, unlike previous figures, we take the negativity as the mean between projections on adjacent qubits rather than the maximum. Figure 16 shows Pearson correlation values of for unmitigated results, and for mitigated results. These values, although lower than predicted, lie within general expectations since higher two-qubit gate error rates should correspond to lower levels of entanglement. Other factors that may impact negativity measurements include relaxation and dephasing time, crosstalk, and single-qubit gate errors.
We also plot the mean device negativity versus mean device CNOT error rate for all IBM Quantum devices in Fig. 16, measuring for unmitigated negativities and for mitigated negativities. The results similarly lie within expectation, indicating the potential utility of our protocol as a scalable whole-device benchmarking tool.
We summarise all graph state experiment results in Table 3. In addition to tabulating device size, quantum volume, and mean device negativity, we display the sizes of the largest connected entanglement graphs with edges above a certain negativity threshold. In detail, columns with label represent the size of the largest connected graph where edges exist only between qubits pairs whose measured negativity is at least of the maximum value. This metric allows us to simultaneously probe the scale and quality of clusters of entanglement. We observe a few general trends. Firstly, the standard deviation in negativities for each device typically decreases once we apply QREM. Similarly, the variance between mean device negativities also diminishes. This may be attributed to the bound on the maximum negativity and variance in mean readout error rates, which range from 1.1% on ibm_lagos to 5.2% on ibmq_quito. Secondly, for devices of similar size, quantum volume is not a good predictor of mean device negativity. For instance, ibm_geneva, which has a quantum volume of 32, has a mean device negativity of (QREM) compared to ibm_kolkata’s , which has a quantum volume of 128. This may be ascribed to a couple of factors, the first being that quantum volume is defined over a subset of qubits, instead of the whole device, and the second being that our entanglement protocol utilises primarily low-depth circuits.
No QREM | |||||||
Device | Qubits | QV | Mean | Whole-Device | |||
lima | 5 | 8 | 5 | 3 | 0 | ✓ | |
belem | 5 | 16 | 5 | 0 | 0 | ✓ | |
quito | 5 | 16 | 5 | 0 | 0 | ✓ | |
manila | 5 | 32 | 5 | 5 | 0 | ✓ | |
jakarta | 7 | 16 | 7 | 0 | 0 | ✓ | |
oslo | 7 | 32 | 7 | 6 | 0 | ✓ | |
nairobi | 7 | 32 | 7 | 3 | 0 | ✓ | |
lagos | 7 | 32 | 7 | 3 | 0 | ✓ | |
perth | 7 | 32 | 7 | 0 | 0 | ✓ | |
guadalupe | 16 | 32 | 16 | 2 | 0 | ✓ | |
toronto | 27 | 32 | 9 | 3 | 0 | ✓ | |
geneva | 27 | 32 | 11 | 2 | 0 | ✗ (26) | |
hanoi | 27 | 64 | 26 | 3 | 0 | ✓ | |
auckland | 27 | 64 | 26 | 13 | 0 | ✓ | |
cairo | 27 | 64 | 27 | 4 | 0 | ✓ | |
mumbai | 27 | 128 | 23 | 4 | 0 | ✓ | |
montreal | 27 | 128 | 8 | 0 | 0 | ✓ | |
kolkata | 27 | 128 | 24 | 9 | 0 | ✓ | |
washington | 127 | 64 | 85 | 6 | 0 | ✗ (121) | |
sherbrooke | 127 | 32 | 125 | 76 | 4 | ✓ | |
brisbane | 127 | - | 125 | 26 | 0 | ✓ | |
seattle | 433 | - | 11 | 3 | 0 | ✗ (184) | |
QREM | |||||||
lima | 5 | 8 | 5 | 5 | 5 | ✓ | |
belem | 5 | 16 | 5 | 5 | 0 | ✓ | |
quito | 5 | 16 | 5 | 5 | 5 | ✓ | |
manila | 5 | 32 | 5 | 5 | 5 | ✓ | |
jakarta | 7 | 16 | 7 | 7 | 7 | ✓ | |
oslo | 7 | 32 | 7 | 7 | 7 | ✓ | |
nairobi | 7 | 32 | 7 | 7 | 7 | ✓ | |
lagos | 7 | 32 | 7 | 7 | 7 | ✓ | |
perth | 7 | 32 | 7 | 7 | 7 | ✓ | |
guadalupe | 16 | 32 | 16 | 16 | 11 | ✓ | |
toronto | 27 | 32 | 27 | 11 | 4 | ✓ | |
geneva | 27 | 32 | 26 | 26 | 25 | ✓ | |
hanoi | 27 | 64 | 27 | 27 | 17 | ✓ | |
auckland | 27 | 64 | 27 | 26 | 13 | ✓ | |
cairo | 27 | 64 | 27 | 27 | 9 | ✓ | |
mumbai | 27 | 128 | 27 | 27 | 23 | ✓ | |
montreal | 27 | 128 | 27 | 24 | 8 | ✓ | |
kolkata | 27 | 128 | 25 | 22 | 11 | ✓ | |
washington | 127 | 64 | 115 | 90 | 43 | ✓ | |
sherbrooke | 127 | 32 | 127 | 127 | 114 | ✓ | |
brisbane | 127 | - | 127 | 124 | 99 | ✓ | |
seattle | 433 | - | 330 | 37 | 11 | ✓ (active qubits) |
Generating 414-Qubit Graph States
Using the same protocol, we characterise bipartite entanglement on a larger 414-qubit graph state prepared on the 433-qubit device , where at the time of experiment, 19 of the 433 device qubits were inoperable. Figure 17 displays the average negativity versus nearest-neighbour qubit pairs. The average qubit pair negativity is found to be 0.115 without QREM, and 0.340 with QREM. We report proportionally higher readout error rates compared to previous devices. The coupling map is displayed in Fig. 18. After mitigating for readout errors, all bipartitions not involving inactive qubits had measured negativities above 0.
V Preserving Whole-Device Graph States Via Dynamical Decoupling
We extend our study of preserving large-scale entanglement on IBM Quantum devices via dynamical decoupling to whole-device graph states. This application holds significant potential as graph states are considered to be more practically relevant than GHZ states on NISQ devices owing to their greater noise resilience. In particular, preserving large-scale entanglement in graph states is crucial for several measurement-based computation schemes, where qubits may experience long idle times due to the distributed manner in which quantum information is processed.
The methodology for testing dynamical decoupling on whole-device graph states parallels the GHZ case. A whole-device graph state is prepared as previously outlined. A variable delay period is inserted between state preparation and whole-device quantum state tomography. We compare negativity decays for circuits without mitigation (free decay), and circuits mitigated with PDD and Hahn echo. The general circuit diagram is shown in Fig. 19 and the PDD circuit in Fig. 6c.
[…]
Whole-Device Graphstate Negativity Decays on a 127-Qubit Device
Time-dependent graphstate experiments are performed on the 127-qubit Eagle r3 backend ibm_brisbane with median µs and µs. The device did not have a published quantum volume at the time of experiment. We vary the delay period from 0 µs up to 12 µs in 1 µs increments for free decays (idle qubits), Hahn echo, and PDD with 4 µs-1 and 8 µs-1 pulse rates (i.e. frequency of X-gates). We implement the double -pulse with a 1:2:1 delay spacing. We perform single sets of experiments for each circuit delay value, executing all circuits with 4096 shots each. We apply readout error mitigation to all results.
Figure 20 shows the negativity over time for each edge in the native-graph state, comparing free decays with various dynamical decoupling configurations. Figure 21 displays the mean device negativity over time, where error bars represent the standard deviation between negativities of individual edges.
Immediately apparent in Fig. 20 are resurgent signals in negativity for several qubit pairs. Furthermore, in Fig. 21, we observe an average increase in device negativity of 0.025 for free decays, and 0.061 for double -pulse, between µs and µs. We also note the sharp negativity oscillations of certain qubit pairs in PDD experiments. While we present data from only a single set of experiments, experiments performed shortly thereafter show similar oscillations in negativity for the same qubit pairs (with some device drift). These oscillations, while initially unexpected, are consistent with signals produced by residual ZZ interactions.
ZZ interactions, also known as ZZ couplings or crosstalk are known to affect weakly anharmonic transmon qubits [32, 33, 34, 35, 36]. ZZ interactions represent a coherent noise process whose effect on idle qubits is effectively a controlled-phase rotation. Therefore, ZZ interactions can generate entanglement between qubit pairs, and conversely, accelerate disentanglement. Assuming no other noise mechanism such as dephasing is present, an isolated ZZ pair interaction produces a signal in the negativity.
With this in mind, we consider the effects of dynamical decoupling sequences on the negativity decays of native-graph states. Figure 21 shows that both PDD experiments demonstrate sizeable improvement in mean entanglement lifetimes over free decay and double -pulse experiments. The double -pulse experiment demonstrates slight improvement over free decays. In addition, increasing the PDD pulse rate does not appear to substantially improve the mean device negativity decay curve. From Fig. 20, we observe that PDD does not completely eliminate revivals in negativity, however does well to prolong the majority of pairwise negativity lifetimes and suppress some coherent noise artifacts.
While we focus on mean improvements in entanglement lifetimes for native graph states, recent results have shown that implementing a precisely timed dynamical decoupling sequence can more effectively cancel the coherent ZZ errors in a 12-qubit ring graph state [61]. A similar approach of tailoring the dynamical decoupling sequence for native graph states may also be beneficial, although can be considerably more complex depending on the scale and connectivity of the underlying graph. For the purposes of improving the mean entanglement lifetime of a large-scale graph state, we show that even a simple PDD sequence incurs significant benefits. Additionally, we note that an adaption of our procedure may be potentially useful in detecting and characterising the coherent noise effects.
VI Discussion
We prepared and studied several GHZ and native-graph states prepared across the range of IBM Quantum devices. In particular, we measured the time-dependent decay of entanglement in these states and verified the efficacy of dynamical decoupling in prolonging entanglement lifetimes.
[…]
For graph states, we developed a bipartite entanglement characterisation protocol that constructs entanglement graphs depicting bipartite entanglement in IBM Quantum devices using as low as a constant 36 circuits. We used the protocol to verify and quantify whole-device bipartite entanglement over 20 different IBM Quantum systems, including three 127-qubit systems. We further showed entanglement across 414 qubits in a 433-qubit device. We then tested dynamical decoupling for preserving qubit pair negativities in a native-graph state prepared on the 127-qubit ibmbrisbane device. We observed coherent noise signals consistent with residual ZZ interactions, which were partially suppressed after application of PDD. PDD led to an overall improvement in mean device bipartite entanglement lifetimes. We also note the potential utility of a running a similar procedure to detect and characterise the coherent noise signals.
Overall, our work highlights both some of the growing capabilities of NISQ devices alongside current limitations through the lens of large-scale entanglement. It also highlights the need for and benefit of noise mitigation and suppression techniques for generating and maintaining large-scale entanglement in NISQ devices.
Acknowledgements.
This work was supported by the University of Melbourne through the establishment of an IBM Quantum Network Hub at the University. We would like to thank Seth Merkel for valuable discussions relating to transmon modelling.References
- Horodecki et al. [2009] R. Horodecki, P. Horodecki, M. Horodecki, and K. Horodecki, Quantum entanglement, Reviews of Modern Physics 81, 865 (2009).
- Einstein et al. [1935] A. Einstein, B. Podolsky, and N. Rosen, Can quantum-mechanical description of physical reality be considered complete?, Physical Review 47, 777 (1935).
- Jozsa and Linden [2003] R. Jozsa and N. Linden, On the role of entanglement in quantum-computational speed-up, Proceedings of the Royal Society of London. Series A: Mathematical, Physical and Engineering Sciences 459, 2011 (2003).
- Vidal [2003] G. Vidal, Efficient classical simulation of slightly entangled quantum computations, Physical Review Letters 91, 10.1103/PhysRevLett.91.147902 (2003).
- Datta et al. [2005] A. Datta, S. T. Flammia, and C. M. Caves, Entanglement and the power of one qubit, Physical Review A 72, 042316 (2005).
- Kendon and Munro [2006] V. M. Kendon and W. J. Munro, Entanglement and its role in Shor’s algorithm, Quantum Information and Computation 6, 10.26421/qic6.7-6 (2006).
- Ekert [1991] A. K. Ekert, Quantum cryptography based on Bell’s theorem, Physical Review Letters 67, 661 (1991).
- Bennett and Wiesner [1992] C. H. Bennett and S. J. Wiesner, Communication via one- and two-particle operators on Einstein-Podolsky- Rosen states, Physical Review Letters 69, 10.1103/PhysRevLett.69.2881 (1992).
- Bennett et al. [1993] C. H. Bennett, G. Brassard, C. Crépeau, R. Jozsa, A. Peres, and W. K. Wootters, Teleporting an unknown quantum state via dual classical and Einstein-Podolsky-Rosen channels, Physical Review Letters 70, 10.1103/PhysRevLett.70.1895 (1993).
- Wootters [1998] W. K. Wootters, Quantum entanglement as a quantifiable resource, Philosophical Transactions of the Royal Society of London. Series A: Mathematical, Physical and Engineering Sciences 356, 1717 (1998).
- Raussendorf et al. [2006] R. Raussendorf, J. Harrington, and K. Goyal, A fault-tolerant one-way quantum computer, Annals of physics 321, 2242 (2006).
- Brun et al. [2006] T. Brun, I. Devetak, and M.-H. Hsieh, Correcting quantum errors with entanglement, science 314, 436 (2006).
- Fowler et al. [2012] A. G. Fowler, M. Mariantoni, J. M. Martinis, and A. N. Cleland, Surface codes: Towards practical large-scale quantum computation, Physical Review A 86, 032324 (2012).
- Chow et al. [2021] J. Chow, O. Dial, and J. Gambetta, IBM Quantum breaks the 100‑qubit processor barrier (2021).
- Madsen et al. [2022] L. S. Madsen, F. Laudenbach, M. F. Askarani, F. Rortais, T. Vincent, J. F. F. Bulmer, F. M. Miatto, L. Neuhaus, L. G. Helt, M. J. Collins, A. E. Lita, T. Gerrits, S. W. Nam, V. D. Vaidya, M. Menotti, I. Dhand, Z. Vernon, N. Quesada, and J. Lavoie, Quantum computational advantage with a programmable photonic processor, Nature 606, 75 (2022).
- Preskill [2018] J. Preskill, Quantum computing in the NISQ era and beyond, Quantum 2, 10.22331/q-2018-08-06-79 (2018).
- Greenberger et al. [1989] D. M. Greenberger, M. A. Horne, and A. Zeilinger, Going Beyond Bell’s Theorem, in Bell’s Theorem, Quantum Theory and Conceptions of the Universe (1989).
- Mooney et al. [2021a] G. J. Mooney, G. A. L. White, C. D. Hill, and L. C. L. Hollenberg, Generation and verification of 27-qubit Greenberger-Horne-Zeilinger states in a superconducting quantum computer, Journal of Physics Communications 5, 095004 (2021a).
- Yang et al. [2022a] B. Yang, R. Raymond, and S. Uno, Efficient quantum readout-error mitigation for sparse measurement outcomes of near-term quantum devices, Physical Review A 106, 10.1103/PhysRevA.106.012423 (2022a).
- Moses et al. [2023] S. A. Moses, C. H. Baldwin, M. S. Allman, R. Ancona, L. Ascarrunz, C. Barnes, J. Bartolotta, B. Bjork, P. Blanchard, and M. Bohn, A race track trapped-ion quantum processor, arXiv preprint arXiv:2305.03828 (2023).
- Wang et al. [2018] Y. Wang, Y. Li, Z. q. Yin, and B. Zeng, 16-qubit IBM universal quantum computer can be fully entangled, npj Quantum Information 4, 10.1038/s41534-018-0095-x (2018).
- Mooney et al. [2019] G. J. Mooney, C. D. Hill, and L. C. L. Hollenberg, Entanglement in a 20-Qubit Superconducting Quantum Computer, Scientific Reports 9, 13465 (2019).
- Mooney et al. [2021b] G. J. Mooney, G. A. White, C. D. Hill, and L. C. Hollenberg, Whole-device entanglement in a 65-qubit superconducting quantum computer, Advanced Quantum Technologies 4, 10.1002/qute.202100061 (2021b).
- Friis et al. [2018] N. Friis, O. Marty, C. Maier, C. Hempel, M. Holzäpfel, P. Jurcevic, M. B. Plenio, M. Huber, C. Roos, R. Blatt, and B. Lanyon, Observation of Entangled States of a Fully Controlled 20-Qubit System, Physical Review X 8, 10.1103/PhysRevX.8.021012 (2018).
- Yang et al. [2022b] B. Yang, R. Raymond, H. Imai, H. Chang, and H. Hiraishi, Testing scalable Bell inequalities for quantum graph states on IBM Quantum devices, IEEE Journal on Emerging and Selected Topics in Circuits and Systems , 1 (2022b).
- Cao et al. [2023] S. Cao, B. Wu, F. Chen, M. Gong, Y. Wu, Y. Ye, C. Zha, H. Qian, C. Ying, S. Guo, Q. Zhu, H. L. Huang, Y. Zhao, S. Li, S. Wang, J. Yu, D. Fan, D. Wu, H. Su, H. Deng, H. Rong, Y. Li, K. Zhang, T. H. Chung, F. Liang, J. Lin, Y. Xu, L. Sun, C. Guo, N. Li, Y. H. Huo, C. Z. Peng, C. Y. Lu, X. Yuan, X. Zhu, and J. W. Pan, Generation of genuine entanglement up to 51 superconducting qubits, Nature 10.1038/s41586-023-06195-1 (2023).
- Kattemölle and van Wezel [2020] J. Kattemölle and J. van Wezel, Conditions for superdecoherence, Quantum 4, 10.22331/q-2020-05-14-265 (2020).
- Monz et al. [2011] T. Monz, P. Schindler, J. T. Barreiro, M. Chwalla, D. Nigg, W. A. Coish, M. Harlander, W. Hänsel, M. Hennrich, and R. Blatt, 14-qubit entanglement: Creation and coherence, Physical Review Letters 106, 10.1103/PhysRevLett.106.130506 (2011).
- Ozaeta and McMahon [2019] A. Ozaeta and P. L. McMahon, Decoherence of up to 8-qubit entangled states in a 16-qubit superconducting quantum processor, Quantum Science and Technology 4, 10.1088/2058-9565/ab13e5 (2019).
- Kaufmann et al. [2017] H. Kaufmann, T. Ruster, C. T. Schmiegelow, M. A. Luda, V. Kaushal, J. Schulz, D. Von Lindenfels, F. Schmidt-Kaler, and U. G. Poschinger, Scalable Creation of Long-Lived Multipartite Entanglement, Physical Review Letters 119, 10.1103/PhysRevLett.119.150503 (2017).
- Wei et al. [2020] K. X. Wei, I. Lauer, S. Srinivasan, N. Sundaresan, D. T. Mcclure, D. Toyli, D. C. Mckay, J. M. Gambetta, and S. Sheldon, Verifying multipartite entangled Greenberger-Horne-Zeilinger states via multiple quantum coherences, Physical Review A 101, 10.1103/PhysRevA.101.032343 (2020).
- McKay et al. [2019] D. C. McKay, S. Sheldon, J. A. Smolin, J. M. Chow, and J. M. Gambetta, Three-Qubit Randomized Benchmarking, Physical Review Letters 122, 10.1103/PhysRevLett.122.200502 (2019).
- Magesan and Gambetta [2020] E. Magesan and J. M. Gambetta, Effective Hamiltonian models of the cross-resonance gate, Physical Review A 101, 52308 (2020).
- Ku et al. [2020] J. Ku, X. Xu, M. Brink, D. C. McKay, J. B. Hertzberg, M. H. Ansari, and B. L. Plourde, Suppression of Unwanted ZZ Interactions in a Hybrid Two-Qubit System, Physical review letters 125, 10.1103/PhysRevLett.125.200504 (2020).
- Sundaresan et al. [2020] N. Sundaresan, I. Lauer, E. Pritchett, E. Magesan, P. Jurcevic, and J. M. Gambetta, Reducing unitary and spectator errors in cross resonance with optimized rotary echoes, PRX Quantum 1, 10.1103/PRXQuantum.1.020318 (2020).
- Zhao et al. [2021] P. Zhao, D. Lan, P. Xu, G. Xue, M. Blank, X. Tan, H. Yu, and Y. Yu, Suppression of Static ZZ Interaction in an All-Transmon Quantum Processor, Physical Review Applied 16, 10.1103/PhysRevApplied.16.024037 (2021).
- Leibfried et al. [2005] D. Leibfried, E. Knill, S. Seidelin, J. Britton, R. B. Blakestad, J. Chiaverini, D. B. Hume, W. M. Itano, J. D. Jost, C. Langer, R. Ozeri, R. Reichle, and D. J. Wineland, Creation of a six-atom ’Schrödinger cat’ state, Nature 438, 10.1038/nature04251 (2005).
- Castelvecchi [2023] D. Castelvecchi, IBM releases first-ever 1,000-qubit quantum chip, Nature 624, 238 (2023).
- Hertzberg et al. [2021] J. B. Hertzberg, E. J. Zhang, S. Rosenblatt, E. Magesan, J. A. Smolin, J. B. Yau, V. P. Adiga, M. Sandberg, M. Brink, J. M. Chow, and J. S. Orcutt, Laser-annealing Josephson junctions for yielding scaled-up superconducting quantum processors, npj Quantum Information 7, 10.1038/s41534-021-00464-5 (2021).
- Takita et al. [2017] M. Takita, A. W. Cross, A. D. Córcoles, J. M. Chow, and J. M. Gambetta, Experimental demonstration of fault-tolerant state preparation with superconducting qubits, Physical Review Letters 119, 10.1103/PhysRevLett.119.180501 (2017).
- Chamberland et al. [2020] C. Chamberland, G. Zhu, T. J. Yoder, J. B. Hertzberg, and A. W. Cross, Topological and subsystem codes on low-Degree graphs with flag qubits, Physical Review X 10, 10.1103/PhysRevX.10.011022 (2020).
- Nation et al. [2021] P. D. Nation, H. Kang, N. Sundaresan, and J. M. Gambetta, Scalable mitigation of measurement errors on quantum computers, PRX Quantum 2, 10.1103/PRXQuantum.2.040326 (2021).
- Smolin et al. [2012] J. A. Smolin, J. M. Gambetta, and G. Smith, Efficient method for computing the maximum-likelihood quantum state from measurements with additive gaussian noise, Physical Review Letters 108, 10.1103/PhysRevLett.108.070502 (2012).
- Viola and Lloyd [1998] L. Viola and S. Lloyd, Dynamical suppression of decoherence in two-state quantum systems, Physical Review A - Atomic, Molecular, and Optical Physics 58, 10.1103/PhysRevA.58.2733 (1998).
- Viola et al. [1999] L. Viola, E. Knill, and S. Lloyd, Dynamical decoupling of open quantum systems, Physical Review Letters 82, 10.1103/PhysRevLett.82.2417 (1999).
- Hahn [1950] E. L. Hahn, Spin echoes, Physical Review 80, 10.1103/PhysRev.80.580 (1950).
- Viola and Knill [2002] L. Viola and E. Knill, Robust dynamical decoupling with bounded controls, Physical Review Letters 90 (2002).
- Khodjasteh and Lidar [2005] K. Khodjasteh and D. A. Lidar, Fault-tolerant quantum dynamical decoupling, Physical Review Letters 95, 10.1103/PhysRevLett.95.180501 (2005).
- Uhrig [2007] G. S. Uhrig, Keeping a quantum bit alive by optimized -pulse sequences, Physical Review Letters 98, 10.1103/PhysRevLett.98.100504 (2007).
- Ezzell et al. [2022] N. Ezzell, B. Pokharel, L. Tewala, G. Quiroz, and D. A. Lidar, Dynamical decoupling for superconducting qubits: a performance survey, arXiv preprint arXiv:2207.03670 (2022).
- Briegel and Raussendorf [2001] H. J. Briegel and R. Raussendorf, Persistent entanglement in arrays of interacting particles, Physical Review Letters 86, 10.1103/PhysRevLett.86.910 (2001).
- Raussendorf and Briegel [2001] R. Raussendorf and H. J. Briegel, A one-way quantum computer, Physical Review Letters 86, 10.1103/PhysRevLett.86.5188 (2001).
- Schlingemann and Werner [2002] D. Schlingemann and R. F. Werner, Quantum error-correcting codes associated with graphs, Physical Review A - Atomic, Molecular, and Optical Physics 65, 10.1103/PhysRevA.65.012308 (2002).
- Markham and Sanders [2008] D. Markham and B. C. Sanders, Graph states for quantum secret sharing, Physical Review A - Atomic, Molecular, and Optical Physics 78, 10.1103/PhysRevA.78.042309 (2008).
- Shettell and Markham [2020] N. Shettell and D. Markham, Graph States as a Resource for Quantum Metrology, Physical Review Letters 124, 10.1103/PhysRevLett.124.110502 (2020).
- Gühne et al. [2005] O. Gühne, G. Tóth, P. Hyllus, and H. J. Briegel, Bell Inequalities for Graph States, Physical Review Letters 95, 120405 (2005).
- Høyer et al. [2006] P. Høyer, M. Mhalla, and S. Perdrix, Resources required for preparing graph states, in Algorithms and Computation, edited by T. Asano (Springer Berlin Heidelberg, Berlin, Heidelberg, 2006) pp. 638–649.
- Aliferis and Leung [2004] P. Aliferis and D. W. Leung, Computation by measurements: A unifying picture, Physical Review A - Atomic, Molecular, and Optical Physics 70, 10.1103/PhysRevA.70.062314 (2004).
- Raussendorf et al. [2003] R. Raussendorf, D. E. Browne, and H. J. Briegel, Measurement-based quantum computation on cluster states, Physical Review A - Atomic, Molecular, and Optical Physics 68, 10.1103/PhysRevA.68.022312 (2003).
- Vidal and Werner [2002] G. Vidal and R. F. Werner, Computable measure of entanglement, Physical Review A - Atomic, Molecular, and Optical Physics 65, 10.1103/PhysRevA.65.032314 (2002).
- Shirizly et al. [2023] L. Shirizly, G. Misguich, and H. Landa, Dissipative Dynamics of Graph-State Stabilizers with Superconducting Qubits, arXiv preprint arXiv:2308.01860 (2023).

### v2

@@
-Generation and Preservation of Large Entangled States on Physical Quantum Devices
+Characterization of entanglement on superconducting quantum computers of up to 414 qubits
 Abstract
@@
 A GHZ fidelity of is measured on a 32-qubit GHZ state, certifying its genuine multipartite entanglement (GME).
-We show a substantial improvement in GHZ decoherence rates for a 7-qubit GHZ state after implementing dynamical decoupling, and observe a linear trend in the decoherence rate of µs for up to qubits, confirming the absence of superdecoherence.
-Additionally, we prepare and characterise fully bipartite entangled native graph states on 22 IBM Quantum devices with qubit counts as high as 414 qubits, all active qubits of the 433-qubit Osprey device.
+We show a substantial improvement in GHZ decoherence rates for a 7-qubit GHZ state after implementing dynamical decoupling, and observe a linear trend in the decoherence rate of µs-1 for up to qubits, confirming the absence of superdecoherence.
+Additionally, we prepare and characterize fully bipartite entangled native graph states on 22 superconducting quantum devices with qubit counts as high as 414 qubits, all active qubits of the 433-qubit IBM Osprey device.
 Analysis of the decay of 2-qubit entanglement within the prepared states shows suppression of coherent noise signals with the implementation of dynamical decoupling techniques.
 Additionally, we observe that the entanglement in some qubit pairs oscillates over time, which is likely caused by residual ZZ-interactions.
-Characterising entanglement in native graph states, along with detecting entanglement oscillations, can be an effective approach to low-level device benchmarking that encapsulates 2-qubit error rates along with additional sources of noise, with possible applications to quantum circuit compilation.
-We develop several tools to automate the preparation and entanglement characterisation of GHZ and graph states.
-In particular, a method to characterise graph state bipartite entanglement using just 36 circuits, constant with respect to the number of qubits.
+Characterizing entanglement in native graph states, along with detecting entanglement oscillations, can be an effective approach to low-level device benchmarking that encapsulates 2-qubit error rates along with additional sources of noise, with possible applications to quantum circuit compilation.
+We develop several tools to automate the preparation and entanglement characterization of GHZ and graph states.
+In particular, a method to characterize graph state bipartite entanglement using just 36 circuits, constant with respect to the number of qubits.
 I Introduction
+| Entangled State | Experiment | Device(s) | Section |
+GHZ
+GHZ state generation and verification up to 32 qubits |
+ibm_washington
+(127 qubits) |
+Section II |
+GHZ
+Extending 7-qubit GHZ state lifetimes with dynamical decoupling |
+ibmq_mumbai
+(27 qubits) |
+Section III.1 |
+GHZ
+Scaling of GHZ decoherence rate versus state size up to 15 qubits |
+ibm_hanoi
+(27 qubits) |
+Section III.2 |
+Graph state
+Scalable whole-device bipartite entanglement characterization up to 414 qubits |
+22 devices,
+see Table 4 |
+Section IV |
+Graph state
+Testing dynamical decoupling for 127-qubit whole-device bipartite entanglement |
+ibm_brisbane
+(127 qubits) |
+Section V |
 Producing large-scale entangled states on many qubits represents an important frontier in the race to scale up physical quantum computers.
@@
 Multipartite entanglement is known for its complex structure [1], and modern quantum devices have passed the number of qubits a classical computer can store an arbitrary quantum state on [14, 15].
-Consequently, as noisy intermediate-scale quantum (NISQ) devices [16] continue to improve in both size and error rates, it becomes paramount to characterise the ability of a quantum computer to generate and maintain large entangled states in a scalable manner.
+Consequently, as noisy intermediate-scale quantum (NISQ) devices [16] continue to improve in both size and error rates, it becomes paramount to characterize the ability of a quantum computer to generate and maintain large entangled states in a scalable manner.
 Verifying multipartite entanglement on a quantum device requires measuring the fidelity of a highly entangled multi-qubit state.
@@
 We observe improvement in mean device entanglement lifetimes after implementing PDD.
+To enhance the paper’s navigability, we provide a guiding table in Table 1 summarizing all main experiments, the devices tested, and the relevant sections of the paper.
 II Generation and Decay of Greenberger-Horne-Zeilinger States
-GHZ states and verifying genuine multipartite entanglement
+II.1 GHZ states and verifying genuine multipartite entanglement
 GHZ states [17] are highly entangled multipartite states that are uniquely fragile to noise, where a single-qubit phase error can destroy the whole-state entanglement.
@@
 Formally, an -qubit GHZ state is expressed as:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
 To certify -qubit GME, it is sufficient to show a GHZ fidelity of greater than 0.5 [37] (note that a GHZ state may be GME with a fidelity under 0.5).
@@
 1.
-Prepare the -qubit GHZ state in the form as exemplified in Fig. 1b.
+Prepare the -qubit GHZ state in the form as exemplified in Fig. 1(b).
 2.
-(Optionally) Apply a refocusing -pulse (i.e. an -gate) on each qubit, producing the state .
+(Optionally) Apply a refocusing -pulse, i.e., an -gate on each qubit.
 3.
@@
 Obtain the overlap signal as the probabilities of measuring the state for different values of .
-The GHZ coherence can then be calculated as where are the overlap signal amplitudes which can be obtained by applying a Fourier transform to :
+The -qubit GHZ coherence can then be calculated as where are the overlap signal amplitudes which can be obtained by applying a Fourier transform to :
 where is the number of angles in the summation.
@@
 Further details including deriving the fidelity from the overlap signal concretely is presented in [31, 18].
-GHZ state embedding on physical topologies
+II.2 GHZ state embedding on physical topologies
 When preparing GHZ states on physical devices, accounting for hardware topology and gate error rates is crucial for maximizing the final state fidelity.
@@
 The method additionally involves selecting qubit subsets with optimized parameters such as low two-qubit gate error rates.
-The exact parameter or combination of parameters to optimise is specified by assigning corresponding weights to the edges of the graph.
+The exact parameter or combination of parameters to optimize is specified by assigning corresponding weights to the edges of the graph.
 We initially considered only the CNOT error rate.
@@
 We note that the algorithm is designed to be compatible with any finite-degree device graph, although its performance on other topologies has not yet been investigated.
-Scalable quantum readout-error mitigation (M3)
+II.3 Scalable quantum readout-error mitigation (M3)
 Readout or measurement errors represent the largest source of noise for low-depth circuits executed on NISQ devices.
@@
 In its simplest form, quantum readout error mitigation (QREM) solves the linear equation
-where is the noisy probability vector returned by the system, is the ideal probability vector in the absence of readout errors (but may include other errors), and is the calibration matrix, where corresponds to the probability of bit string being measured as bit string .
+where is the noisy probability vector returned by the system, is the ideal probability vector in the absence of readout errors (but may include other errors), and is the calibration matrix.
 There are limitations to QREM, with the most significant one being the exponential scaling of classical resources required to solve for with respect to .
@@
 M3 relies on two main optimizations: subspace reduction of the full calibration matrix based on noisy input bit strings, and implementation of a matrix-free, iterative method for solving the system of linear equations.
-The method involves directly computing matrix elements as
+For two bit-strings , where qubit 0 corresponds to the least significant bit, the method directly computes matrix elements as
+where corresponds to the probability of the -th qubit being in state and measured in state .
 This error mitigation technique has some caveats.
@@
 We apply the correction algorithm with all the default settings, which include correcting bit strings up to a Hamming distance equal to GHZ size , a convergence tolerance of the iterative method of , and a maximum number of iterations of 25.
-Verifying 32-qubit GHZ states on a 127-qubit Eagle Processor
+II.4 Verifying 32-qubit GHZ states on a 127-qubit Eagle Processor
 We prepare and measure the fidelities of tree-type GHZ states of sizes , on the 127-qubit ibm_washington device.
@@
 III Preserving GHZ States via Dynamical Decoupling
-@C=0.8em @R=0.7em
-[…]
-[…]
-[…]
-[…]
-@C=0.8em @R=0.5em
-[…]
-[…]
-[…]
-[…]
-@C=0.8em @R=0.85em
-[…]
-[…]
-[…]
-[…]
-[…]
+III.1 Hahn Echo and Periodic Dynamical Decoupling for GHZ States
 In this section, we explore the potential of dynamical decoupling techniques in prolonging the lifetimes of GHZ states on IBM Quantum devices.
@@
 6.
-Hahn Echo and Periodic Dynamical Decoupling for 7-Qubit GHZ States
-Decay experiments are performed for 7-qubit GHZ states prepared on the ibmq_mumbai device consisting of 27 qubits and a quantum volume of 128.
+Decay experiments are performed for 7-qubit GHZ states prepared on the ibmq_mumbai device consisting of 27 qubits and a quantum volume [50] of 128.
 For free decays, we increment the circuit delay by 1 µs up to a maximum delay of 15 µs.
@@
 As shown in Fig.
-7a, neither Hahn echo nor PDD led to a marked improvement of GHZ population times.
+7(a), neither Hahn echo nor PDD led to a marked improvement of GHZ population times.
 In fact, PDD appears to accelerate the decay of GHZ populations.
@@
 In fact, as shown in the PDD curve, repeated application of -gates only introduces additional noise from single-qubit gate errors.
-Figure 7b shows substantial improvement in GHZ coherence times in experiments with Hahn echo and PDD.
+Figure 7(b) shows substantial improvement in GHZ coherence times in experiments with Hahn echo and PDD.
 The GHZ coherences quantify the non-classical correlations between the and states.
-For free decays, this correlation drops to effectively zero by around µs.
+For free decays, this correlation drops to approximately 0.1 by µs.
 In contrast, the PDD curve maintains a measured GHZ coherence of after µs.
@@
 The decay of GHZ state populations and coherences are combined in Fig.
-7c, which plots the GHZ state fidelities as a function of circuit delay .
+7(c), which plots the GHZ state fidelities as a function of circuit delay .
 Both Hahn echo and PDD are shown to be effective techniques for preserving GHZ fidelities.
@@
 Next, we evaluate the scaling of GHZ decoherence rates as a function of state size.
-Scaling of Decoherence Rates and GHZ Size
+III.2 Scaling of Decoherence Rates and GHZ Size
 Studying the decoherence rates of multipartite states as a function of their size may provide vital insight into the noise underlying a quantum system.
@@
 Superdecoherence describes the coupling of qubits to a single reservoir, which cause qubit decoherence rates to scale with the size of the system [27].
-Such an effect is detrimental to the realisation of large-scale, fault-tolerant quantum computers.
+Such an effect is detrimental to the realization of large-scale, fault-tolerant quantum computers.
 GHZ states are particularly convenient for detecting superdecoherence due to their high sensitivity to noise.
@@
 the system experiences superdecoherence), we expect GHZ decoherence rates to scale polynomially.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
 GHZ decoherence rates as a function of state size have been studied on ion-trap and superconducting quantum devices for up to state size 6 and 8, respectively [28, 29].
@@
 where is the initial coherence, and is the decoherence rate, where its reciprocal is the GHZ coherence time.
-GHZ Decoherence Scaling Results
-Figure 9a plots the -qubit GHZ state coherences (normalized) as a function of circuit delay for on ibm_hanoi, which consists of 27 qubits and a quantum volume of 64.
+Figure 9(a) plots the -qubit GHZ state coherences (normalized) as a function of circuit delay for on ibm_hanoi, which consists of 27 qubits and a quantum volume of 64.
 We increase total delay in increments of 2.5 µs.
@@
 In order to quantify this trend, we plot the decoherence rate and the GHZ size , shown in Fig.
-9b.
+9(b).
 We take the error from the standard deviation, which we derive from the covariance matrix produced by the fitting algorithm.
-In agreement with Ozaeta & McMahon, we are able to well fit the data with the linear trendline µs with —now extending up to .
+In agreement with Ozaeta & McMahon, we are able to well fit the data with the linear trendline µs-1 with —now extending up to .
 (With an anomoalous data point at ).
@@
 | [] (MQC) | [µs] (MQC) | [µs] [29] | |
-1 | - | - | |
-2 | - | - | |
-3 | |||
-4 | - | - | |
-5 | |||
-6 | - | - | |
-7 | |||
-8 | - | - | |
-9 | - | ||
-11 | - | ||
-13 | - | ||
-15 | - |
-We summarise the GHZ coherence times on ibm_hanoi in Table 1.
+| 1 | - | - | |
+| 2 | - | - | |
+| 3 | |||
+| 4 | - | - | |
+| 5 | |||
+| 6 | - | - | |
+| 7 | |||
+| 8 | - | - | |
+| 9 | - | ||
+| 11 | - | ||
+| 13 | - | ||
+| 15 | - |
+We summarize the GHZ coherence times on ibm_hanoi in Table 2.
 For easy comparison, we include the measured GHZ coherence times from Ozaeta & McMahon’s experiments.
@@
 The improvement in GHZ coherence times may be attributed to hard improvements in combination with superior coherence detection methods, incorporating dynamical-decoupling based techniques.
-IV Efficient Bipartite Entanglement Characterisation in Whole-Device Graph States
-Graph States
+IV Efficient Bipartite Entanglement Characterization in Whole-Device Graph States
+IV.1 Graph States
 Graph states are a class of entangled multi-qubit states that are defined with respect to a connected graph.
-They are a generalisation of cluster states [51], and hence form a universal basis for measurement-based computation [52].
+They are a generalization of cluster states [52], and hence form a universal basis for measurement-based computation [53].
 Graph states are additionally useful for quantum error correcting codes [54], quantum secure communication [55], quantum metrology [56], and probing Bell inequalities over multi-partite systems [57].
@@
 where , is the set of edges connecting graph containing vertices (qubits), and represents a controlled-phase gate between adjacent qubits and .
-We can equivalently define the stabilising operator for each qubit in :
+We can equivalently define the stabilizing operator for each qubit in :
 where are Pauli operators acting on qubit and is the set of qubits adjacent to . Thus is the simultaneous eigenstate of operators following .
@@
 The qubit layout is represented in Fig.
-10a where nodes represent qubits and edges display possible CNOT operations.
+10(a) where nodes represent qubits and edges display possible CNOT operations.
 Since the graph has a maximum node degree of three, its graph state preparation circuit has a two-qubit gate depth of three.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-To characterise bipartite entanglement across an entire device, we prepare a native-graph state containing every edge.
+To characterize bipartite entanglement across an entire device, we prepare a native-graph state containing every edge.
 We then perform full quantum state tomography on every locally entangled bipartition corresponding to each edge on the device.
@@
 A maximally entangled Bell state has a negativity of , whereas a fully separable state has a negativity of .
-Although there are many other entanglement measures, such as entropy of entanglement and concurrence, negativity is chosen because it is easy to compute and is related to the teleportation distance [60].
-Specifically, the minimum teleportation distance achievable with state , acting on , is bound by
+Although there are many other entanglement measures, including more complex multipartite entanglement witnesses, negativity is an entanglement monotone that is simple to compute.
+A non-zero measurement for the negativity on a 2-qubit state is a necessary and sufficient condition for entanglement.
+This makes it a great choice for bipartite entanglement on graph states, where edges can be reduced to Bell states.
+Additionally, negativity is related to the minimum teleportation distance achievable with state acting on :
 A device is said to be whole-device entangled if every qubit is connected to the main graph where edges correspond to qubit pairs with a measured negativity of .
 This is distinct from saying the qubits are genuinely multipartite entangled (which follows a more strict criteria), but rather, there exist no bipartition of qubits on the device that results in separable states.
-Bipartite Entanglement Characterisation Protocol
-We develop a protocol to efficiently characterise bipartite entanglement on quantum computers, inspired by experiments from Mooney et al.
+IV.2 Bipartite Entanglement Characterization Protocol
+We develop a protocol to efficiently characterize bipartite entanglement on quantum computers, inspired by experiments from Mooney et al.
 [22, 23].
-Development of the scheme is driven by three main design principles: to devise an entanglement characterisation protocol that is highly automated, scalable, and architecture-independent.
-We implement the protocol in Python, utilising the Qiskit API to interface with IBM superconducting quantum devices.
+Development of the scheme is driven by three main design principles: to devise an entanglement characterization protocol that is highly automated, scalable, and architecture-independent.
+We implement the protocol in Python, utilizing the Qiskit API to interface with IBM superconducting quantum devices.
 However, the techniques and procedures in the program are generally applicable.
@@
 The first and second component, which contain techniques unique to this work [22, 23], are elaborated below.
-IV.0.1 Native-Graph State Preparation
+IV.2.1 Native-Graph State Preparation
 Similar to the GHZ case, the objective is to embed depth-optimal circuits using an automated routine.
@@
 Additional graph state embeddings on other IBM physical layouts are shown in the results section below.
-IV.0.2 Parallel Quantum State Tomography
+IV.2.2 Parallel Quantum State Tomography
 In previous graph state experiments [22, 23], full quantum state tomography (QST) is performed on one qubit pair (and its neighbours) at a time.
-Using such method, the number of circuits required to fully characterise a device increases linearly with the number of edges.
+Using such method, the number of circuits required to fully characterize a device increases linearly with the number of edges.
 Our procedure improves upon this by executing QST in parallel.
 In detail, we perform simultaneous basis measurements on non-overlapping sets of qubits, where each set defines a target qubit pair and its neighbours.
-By grouping sets into batches and performing parallel QST for batches at a time, we can fully characterise bipartite entanglement on any size device (provided invariable qubit topology) with a constant number of circuits.
-Using a specialised scheduling scheme for the heavy hex hardware layout, the number of required batches can be reduced to only four when allowing Bell pairs to share neighbours.
+By grouping sets into batches and performing parallel QST for batches at a time, we can fully characterize bipartite entanglement on any size device (provided invariable qubit topology) with a constant number of circuits.
+Using a specialized scheduling scheme for the heavy hex hardware layout, the number of required batches can be reduced to only four when allowing Bell pairs to share neighbours.
 This can be achieved by using an alternate tiling of two unit cells (able to be rotated 180°), where each unit cell is composed of six edges and each edge is assigned to one of four batches.
@@
 In our implementation, we prevent target pairs from sharing neighbours for practical convenience.
-Table 2 lists the number of batches required to characterise IBM Quantum devices up to 433 qubits in size.
+Table 3 lists the number of batches required to characterize IBM Quantum devices up to 433 qubits in size.
 | Device(s) | Qubits | Batches | Circuits |
@@
 Nevertheless, the procedure’s main utility lies in reducing the number of tomography circuits to a roughly constant number in addition to being compatible with various qubit topologies.
-Bipartite Entanglement on IBM Quantum Devices
-We characterise bipartite entanglement on all IBM Quantum devices accessible by the University of Melbourne IBM Quantum Hub.
+IV.3 Bipartite Entanglement on IBM Quantum Devices
+We characterize bipartite entanglement on all IBM Quantum devices accessible by the University of Melbourne IBM Quantum Hub.
 At the time of experiment, these include four 5-qubit systems, five 7-qubit systems, one 16-qubit system, eight 27-qubit systems, three 127-qubit systems, and one 433-qubit system, totaling to 22 systems.
@@
 The results similarly lie within expectation, indicating the potential utility of our protocol as a scalable whole-device benchmarking tool.
-We summarise all graph state experiment results in Table 3.
+We summarize all graph state experiment results in Table 4.
 In addition to tabulating device size, quantum volume, and mean device negativity, we display the sizes of the largest connected entanglement graphs with edges above a certain negativity threshold.
@@
 For instance, ibm_geneva, which has a quantum volume of 32, has a mean device negativity of (QREM) compared to ibm_kolkata’s , which has a quantum volume of 128.
-This may be ascribed to a couple of factors, the first being that quantum volume is defined over a subset of qubits, instead of the whole device, and the second being that our entanglement protocol utilises primarily low-depth circuits.
+This may be ascribed to a couple of factors, the first being that quantum volume is defined over a subset of qubits, instead of the whole device, and the second being that our entanglement protocol utilizes primarily low-depth circuits.
 | No QREM | |||||||
@@
 | brisbane | 127 | - | 127 | 124 | 99 | ✓ | |
-seattle | 433 | - | 330 | 37 | 11 | ✓ (active qubits) |
-Generating 414-Qubit Graph States
-Using the same protocol, we characterise bipartite entanglement on a larger 414-qubit graph state prepared on the 433-qubit device , where at the time of experiment, 19 of the 433 device qubits were inoperable.
+| seattle | 433 | - | 330 | 37 | 11 | ✓ (active qubits) | |
+IV.4 Generating 414-Qubit Graph States
+Using the same protocol, we characterize bipartite entanglement on a larger 414-qubit graph state prepared on the 433-qubit device , where at the time of experiment, 19 of the 433 device qubits were inoperable.
 Figure 17 displays the average negativity versus nearest-neighbour qubit pairs.
@@
 19 and the PDD circuit in Fig.
-6c.
-[…]
-[…]
-[…]
-[…]
-[…]
-Whole-Device Graphstate Negativity Decays on a 127-Qubit Device
+6(b).
+V.1 Whole-Device Graphstate Negativity Decays on a 127-Qubit Device
 Time-dependent graphstate experiments are performed on the 127-qubit Eagle r3 backend ibm_brisbane with median µs and µs.
@@
 For the purposes of improving the mean entanglement lifetime of a large-scale graph state, we show that even a simple PDD sequence incurs significant benefits.
-Additionally, we note that an adaption of our procedure may be potentially useful in detecting and characterising the coherent noise effects.
+Additionally, we note that an adaption of our procedure may be potentially useful in detecting and characterizing the coherent noise effects.
 VI Discussion
@@
 […]
-For graph states, we developed a bipartite entanglement characterisation protocol that constructs entanglement graphs depicting bipartite entanglement in IBM Quantum devices using as low as a constant 36 circuits.
+For graph states, we developed a bipartite entanglement characterization protocol that constructs entanglement graphs depicting bipartite entanglement in IBM Quantum devices using as low as a constant 36 circuits.
 We used the protocol to verify and quantify whole-device bipartite entanglement over 20 different IBM Quantum systems, including three 127-qubit systems.
@@
 PDD led to an overall improvement in mean device bipartite entanglement lifetimes.
-We also note the potential utility of a running a similar procedure to detect and characterise the coherent noise signals.
+We also note the potential utility of a running a similar procedure to detect and characterize the coherent noise signals.
 Overall, our work highlights both some of the growing capabilities of NISQ devices alongside current limitations through the lens of large-scale entanglement.
@@
 - Uhrig [2007] G. S. Uhrig, Keeping a quantum bit alive by optimized -pulse sequences, Physical Review Letters 98, 10.1103/PhysRevLett.98.100504 (2007).
+- Cross et al.
+[2019] A.
+W.
+Cross, L.
+S.
+Bishop, S.
+Sheldon, P.
+D.
+Nation, and J.
+M.
+Gambetta, Validating quantum computers using randomized model circuits, Physical Review A 100, 10.1103/PhysRevA.100.032328 (2019).
 - Ezzell et al. [2022] N. Ezzell, B. Pokharel, L. Tewala, G. Quiroz, and D. A. Lidar, Dynamical decoupling for superconducting qubits: a performance survey, arXiv preprint arXiv:2207.03670 (2022).